OpenStudy (anonymous):

if a,b,c are real and (bc + ca + ab)^2 > k (a+b+c) , then k =?

OpenStudy (anonymous):

@satellite73 @experimentX @ujjwal @UnkleRhaukus @Ishaan94 @dumbcow @shubhamsrg

OpenStudy (experimentx):

k=0

OpenStudy (anonymous):

lol....\(k=0\) workin

OpenStudy (shubhamsrg):

LHS = b^2c^2 + c^2a^2 + a^2b^2 + 2(a^2bc + ab^2c + abc^2) =b^2c^2 + c^2a^2 + a^2b^2 + 2abc(a+b+c) hence, (ab + bc +ca)^2 = (some positive no.) + 2abc(a+b+c) thus (ab+bc+ca)^2 > 2abc(a+b+c) then i guess k =2abc ?? note that (ab+bc+ca)^2 = 2abc(a+b+c) when a=b=c

OpenStudy (anonymous):

The answer should be 3

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (shubhamsrg):

ohh..lol..

OpenStudy (shubhamsrg):

leme try and see..

OpenStudy (anonymous):

ok

OpenStudy (shubhamsrg):

i dont see 3 as an ans.. for a=0,b=1,c=0 ,,3 fails!!

OpenStudy (shubhamsrg):

2abc doesnt seem to fail for me ,,,

OpenStudy (shubhamsrg):

if we see it as AM>GM , (a+b)^2 >4ab =>a^2 + b^2 >2ab similarily, b^2 + c^2 > 2bc c^2 + a^2 > 2ca adding all, a^2 + b^2 +c^2 > ab+bc+ca a^2 + b^2 + c^2 + 2(ab+bc+ca) > 2(ab+bc+ca) + ab+bc+ca (a+b+c)^2 > 3(ab+bc+ca) lol..it turned the other way around this time!! :D

OpenStudy (shubhamsrg):

its symmetrically different from what the ques asks..

OpenStudy (anonymous):

@Yahoo! plz check the question again..........lool

OpenStudy (experimentx):

well ... i would say the same!! if k = 3